# Meta-Learning for Fast Adaptation in Intent Inferral on a Robotic Hand Orthosis for Stroke

### v1

@@ I INTRODUCTION
In upper-limb devices, user-driven interfaces often employ a model that predicts user intent from surface electromyography (EMG) data. Such a model needs to not only capture the relationship between signal and intent across individuals, but also be robust to some amount of variation from factors like user fatigue and sensor migration. More importantly, stroke survivors’ hand spasticity and muscle tone can change drastically across different sessions on different days, commonly known to the community as intersession concept drift. A session is defined to be a single use of the device between donning and doffing the device. These challenges are common to most applications of EMG-driven robotic systems but are especially difficult to mitigate in the stroke rehabilitation space.
Data scarcity in the stroke survivor demographic is an additional key aspect. Both collecting and annotating data are burdensome for therapists and stroke survivors, and few open-source datasets currently exist. This limits our abilities to address the challenges discussed above by simply learning high-capacity intent inferral models trained on very large amounts of data. While more data always helps, we believe that the field also stands to benefit from the investigation of alternative, complementary approaches.
@@
In this paper, we propose using a meta-learning (or learning to learn) approach to training intent inferral models on a robotic orthosis. With this approach, a model can be explicitly optimized to adapt to new environments and tasks when labeled training examples are limited. In the context of wearable robotics, this means training intent inferral models that can quickly adapt to new subjects and new sessions with as little subject- or session-specific data as possible. Our directional goal is to train models that are as effective as possible in making use of the limited data available for any new patient or session.
Specifically, we leverage real clinical data collected from different subjects and different orthosis use conditions in order to train a base model for intent inferral using Model-Agnostic Meta-Learning (MAML) [1]. The MAML training procedure explicitly rewards a base model that is easily adaptable when presented with new data. When the intent inferral model must be used for a new patient or on a new day, we only need a small set of new, session-specific training data to fine-tune the base model to the new condition. We find that our method, which we dub MetaEMG, is particularly effective at learning good intent classifiers with only a few training epochs, and that classifiers trained via MetaEMG are better at adapting to new subjects. In summary, our main contributions are as follows:
@@ III-C MetaEMG
Under our task formulation, the proposed MetaEMG is an application of the Model-Agnostic Meta-Learning (MAML) algorithm to EMG intent inferral for stroke. MetaEMG aims to find initial model parameters that best allow it to adapt to a new session or subject from a limited dataset with only a few fine-tuning epochs. In MetaEMG, two nested optimization loops are used to fine-tune a base model for different tasks using small amounts of data, as shown in Fig. 3.
We train models in a way that directly mirrors the use case of our intent inferral models. In each inner loop of MetaEMG, a base model is fine-tuned on the support set $\mathcal{D}_{s}$ of each training task, and its ability to adapt to the task is measured by calculating the query loss $\mathcal{L}_{\mathcal{D}_{q}}$ of that task. Importantly, the fine-tuning step of the inner loop is performed via gradient descent on $\mathcal{L}_{\mathcal{D}_{s}}$ while the base model update of the outer loop is performed via gradient descent on $\mathcal{L}_{\mathcal{D}_{q}}$ with second-order derivatives. By updating the base model using $\mathcal{L}_{\mathcal{D}_{q}}$, the training process primarily favors base model updates that allow it to generalize to the query set $\mathcal{D}_{q}$. For EMG intent inferral, we hypothesize that this training structure can produce base models that can better learn from a small subject or session-specific dataset with only a few fine-tuning epochs.
@@ IV Data Collection and Session Conditions
All data used in this study was collected using MyHand, a robotic exotendon device comprised of a system of motors and linkages which assist the user in opening and closing their hand, as shown in Fig. 1. The device rests on the user’s forearm and through a series of tendons connected to the user’s fingertips assists the hand in opening and closing a grasp. EMG sensor data is recorded from from an 8-electrode armband (Myo, Thalmic Labs). Once a command is issued to either open or close the hand (for example, by an intent inferral model as the one that is the focus of this study), the orthosis engages the motor and either retracts or extends the exotendons on the dorsal side of the finger. Finger extension is actively assisted by the device through tendon retraction, while finger flexion is allowed unimpeded via the subject’s own strength. Additional hardware details can be found in our previous studies [20, 21].
@@ IV-A Data
Critically, the fact that we collect our data using a complete, functional assistive device means that our collection protocol can span the conditions that are meaningful for real-life operations. Our previous work [3, 2] has shown that, in our patient population, EMG signals display significant differences when collected in isolation (i.e. using just an armband and no other hardware) vs. when collected in conjunction with an active orthosis providing assistance. Thus, in our method, we collect data in four different experimental conditions as follows:
@@ IV-B Pre-processing
The raw output of the EMG sensor is an 8-channel time-series signal (shown in Fig. 1) sampled at 100 Hz which is pre-processed before going into our intent inferral model. The labels attached to each sample are the verbal cues given to the subject during the session, and in a 2-second time interval, there are 200, 8-channel samples. During pre-processing, the 8-channel EMG signal is clipped to the range $[0,1000]$ to discard any outliers, and each channel is re-scaled to the range $[-1,1]$. Finally, the scaled 8-channel signal is binned into 2 second-long 8-channel data windows, following a sliding window stride length of 10ms.
@@ V Experimental Setup
We evaluate MetaEMG in a series of experiments conducted on data collected using the MyHand orthosis and involving five chronic stroke survivors of moderate muscle tone (MAS $\leq 2$) and varying degrees of arm impairment (FM-UE scores listed in Table I) in sessions approved by the Columbia University Institutional Review Board (IRB-AAAS8104).
@@ VI Results & Discussion
TABLE II: Session adaptation results. Classification accuracy and standard deviation across 5 stroke subjects. Results are averages across 3 separate randomly generated training seeds. The best results for each subject are reported in bold.
@@
TABLE III: Subject adaptation results. Classification accuracy and standard deviation across 5 stroke subjects. Results are averages across 3 separate randomly generated training seeds. The best results for each subject are reported in bold.

### v2

@@ I INTRODUCTION
In upper-limb devices, user-driven interfaces often employ a model that predicts user intent from surface electromyography (EMG) data. Such a model needs to not only capture the relationship between signal and intent across individuals, but it also needs to be robust to some amount of variation from factors like user fatigue and sensor migration. In the stroke survivor demographic, hand spasticity and muscle tone can also change drastically across different sessions (single use of the device between donning and doffing) on different days (commonly known to the community as intersession concept drift), adding an additional challenge for model development in the stroke rehabilitation space.
While utilizing complex models to have a better chance of generalization across the large variation in data is a promising research direction, data scarcity in the stroke survivor demographic is a major limiting factor. Both collecting and annotating data are burdensome for therapists and stroke survivors, and very few open-source datasets currently exist. This restricts our ability to address the challenges discussed above by simply learning high-capacity intent inferral models trained on large amounts of data. While more data always helps, we believe that the field also stands to benefit from the investigation of alternative, complementary approaches that enable us to to train larger models with potentially less data.
@@
In this paper, we propose a meta-learning (or learning to learn) approach to training intent inferral models on a robotic orthosis. With this approach, a model can be explicitly optimized to adapt to new environments and tasks when labeled training examples are limited. In the context of wearable robotics, this means training intent inferral models that can quickly adapt to new subjects and new sessions with as little subject- or session-specific data as possible. Our directional goal is to train models that are as effective as possible in making use of the limited data available for any new patient or session.
Specifically, we leverage real clinical data collected from different subjects and different orthosis use conditions in order to train a base model for intent inferral using Model-Agnostic Meta-Learning (MAML) [1]. The MAML algorithm explicitly rewards a base model that is easily adaptable when presented with new data. When the intent inferral model must be used for a new patient or on a new day, we only need a small set of new, session-specific training data to fine-tune the base model to the new condition. We find that our method, which we dub MetaEMG, is particularly effective at learning good intent classifiers with only a few training epochs, and that classifiers trained via MetaEMG are better at adapting to new subjects. In summary, our main contributions are as follows:
@@ III-C MetaEMG
Under our task formulation, the proposed MetaEMG is an application of the Model-Agnostic Meta-Learning (MAML) [1] algorithm to EMG intent inferral. MAML is a meta learning algorithm that focuses on developing models with enhanced adaptability. Rather than optimizing a model for a specific task, MAML aims to create a model that can quickly adapt to new tasks with minimal additional training essentially training the model to be an efficient learner. By doing so, this algorithm can produce models that perform well across a diverse range of tasks, even when provided with only a small number of examples for each new task [1].
MetaEMG aims to find the optimal initial model parameters for adapting a base model to a new session or subject from a limited dataset in only a few fine-tuning epochs. In MetaEMG, two nested optimization loops are used to fine-tune a base model for different tasks using small amounts of data, as shown in Fig. 3 and Alg. 1.
Algorithm 1 MAML for EMG intent inferral
1:$T$: collection of NN tasks $\tau_{i}$, where each $\tau_{i}$ is a single session recording
[…]
3:$M$, $K$: epochs for inner and outer optimization loops
4:randomly initialize base model parameters $\theta$
[…]
6:while  $k\leq K$  do (outer loop)
7:  for each task $\tau_{i}$ do
8:   $\hat{\theta_{i}}\leftarrow\theta$
9:   extract support and query sets $\mathcal{D}_{s}^{i}$ and $\mathcal{D}_{q}^{i}$
10:   while  $m\leq M$ do (inner loop)
[…]
12:   end while
[…]
15:  end for
[…]
17:end while
One of the advantages of formulating intent inferral in the meta learning framework is that it allows us to train models in a way that directly mirrors their use case. In the inner loop of MetaEMG, a base model is fine-tuned on the support set $\mathcal{D}_{s}$ of a training task, and that base model’s ability to adapt to the task is measured through the query loss $\mathcal{L}_{\mathcal{D}_{q}}$ of that task (Alg. 1, steps 7 - 10). The query loss is simply the cross-entropy loss for the query set of the task computed between the predictions and ground-truth labels. This process simulates the real-life use case of fine-tuning an intent classifier on a small amount of demonstration data and using that model for real-time inference in a session with the device.
Moreover, the base model update is performed once per iteration of the outer optimization loop. The query loss is used to accumulate the gradients on the base model parameters across all of the training tasks until a backward step is performed once all of those training tasks have been seen (Alg. 1, steps 11 - 13).
Importantly, the fine-tuning in the inner loop is performed via gradient descent on $\mathcal{L}_{\mathcal{D}_{s}}$ (taken with respect to the latest parameter set $\hat{\theta_{i}}$), while the base model update of the outer loop is performed via gradient descent on $\mathcal{L}_{\mathcal{D}_{q}}$ (taken with respect to the base model parameter set $\theta$) which requires backpropagation through the inner optimization and thus higher-order gradients. By updating the base model using $\mathcal{L}_{\mathcal{D}_{q}}$, the training process favors base model updates that allow it to on average generalize to both the support $\mathcal{D}_{s}$ and query $\mathcal{D}_{s}$ sets of the training tasks despite only being fine-tuned on the support sets. For EMG intent inferral, we hypothesize that this training structure can produce intent classifiers that are less likely to overfit when fine-tuned to small amounts of demonstration data.
@@ IV Data Collection and Session Conditions
All data used in this study was collected using MyHand, a robotic exotendon device comprised of a system of motors and linkages which assist the user in opening and closing their hand, as shown in Fig. 1. The device rests on the user’s forearm and through a series of tendons connected to the user’s fingertips assists the hand in opening and closing a grasp. EMG sensor data is recorded from an 8-electrode armband (Myo, Thalmic Labs). Once a command is issued to either open or close the hand (for example, by an intent inferral model as the one that is the focus of this study), the orthosis engages the motor and either retracts or extends the exotendons on the dorsal side of the finger. Finger extension is actively assisted by the device through tendon retraction, while finger flexion is allowed unimpeded via the subject’s own strength. Additional hardware details can be found in our previous studies [20, 21].
@@ IV-A Data
Critically, the fact that we collect our data using a complete, functional assistive device means that our collection protocol can span the conditions that are meaningful for real-life operations. Our previous work [3, 2] has shown that, in our patient population, EMG signals display significant differences when collected in isolation (i.e. using just an armband and no other hardware) vs. when collected in conjunction with an active orthosis assisting finger movement. Thus, in our method, we collect data in four different experimental conditions as follows:
@@ IV-B Pre-processing
The raw output of the EMG sensor is an 8-channel time-series signal (shown in Fig. 1) sampled at 100 Hz which is pre-processed before going into our intent inferral model. The labels attached to each sample are the verbal cues given to the subject during the session, and in a 2-second time interval, there are 200, 8-channel samples. During pre-processing, the 8-channel EMG signal is clipped to the range $[0,1000]$ to discard any outliers, and each channel is re-scaled to the range [1,1][-1,1]. Finally, the scaled 8-channel signal is binned into 2 second-long 8-channel data windows, following a sliding window stride length of 10ms. While muscle activation is known to occur in a much smaller time window, we hypothesize that 2 seconds encapsulates the total reaction time from interpreting the verbal command to actual movement exertion.
@@ V Experimental Setup
We evaluate MetaEMG in a series of experiments conducted on data collected using the MyHand orthosis and involving five chronic stroke survivors of moderate muscle tone (MAS $\leq 2$) and varying degrees of arm impairment (FM-UE scores listed in Table I) in sessions approved by the Columbia University Institutional Review Board (IRB-AAAS8104). Subjects are generally able to actively close their hands but not open them, which is one of the motivating factors in both the robotic orthosis design as well as the intent inferral structure.
@@ VI Results & Discussion
TABLE II: Session adaptation results. Classification accuracy and standard deviation across 5 stroke subjects. Results are averaged across 3 separate randomly generated training seeds. The best results for each subject are reported in bold.
@@
TABLE III: Subject adaptation results. Classification accuracy and standard deviation across 5 stroke subjects. Results are averaged across 3 separate randomly generated training seeds. The best results for each subject are reported in bold.